# Exploratory study on the masses of odd-$Z$ nuclei and $r$-process simulation based on the deformed relativistic Hartree-Bogoliubov theory in continuum

### v1

@@ II.2 Classical $r$-process
In the astrophysical environments with high-temperature $T\gtrsim 1.5~{}\mathrm{GK}$ and high neutron density $n_{n}\gtrsim 10^{20}\mathrm{~{}cm}^{-3}$, the equilibrium between neutron capture and photodisintegration reactions can be achieved, and the abundance ratios of neighboring isotopes on an isotopic chain can be obtained by the Saha equation [90, 91, 92]
$\frac{Y(Z,A+1)}{Y(Z,A)}=n_{n}\left(\frac{2\pi\hbar^{2}}{m_{\mu}kT}\right)^{3/2}\frac{G(Z,A+1)}{2G(Z,A)}\left(\frac{A+1}{A}\right)^{3/2}\times\exp\left[\frac{S_{n}(Z,A+1)}{kT}\right],$ (10)
where $Y(Z,A),S_{n}(Z,A)$, and $G(Z,A)$ are, respectively, the abundance, one-neutron separation energy, and partition function of nuclide $(Z,A)$, and $\hbar,k$, and $m_{\mu}$ are the Planck constant, Boltzmann constant, and atomic mass unit, respectively. Note that the neutron separation energy $S_{n}$ deduced from nuclear masses appears in the exponential, suggesting the importance of nuclear masses in the equilibrium.
@@
$\frac{dY(Z)}{dt}=Y(Z-1)\sum_{A}P(Z-1,A)\lambda_{\beta}^{Z-1,A}-Y(Z)\sum_{A}P(Z,A)\lambda_{\beta}^{Z,A},$ (11)
where $\lambda_{\beta}^{Z,A}$ is the $\beta$ decay rate of the nucleus $(Z,A)$, $Y(Z)=\sum_{A}Y(Z$, $A)=\sum_{A}P(Z,A)Y(Z)$ is the total abundance of each isotopic chain. By using Eqs. (10) and (11), the abundance of each isotope can be determined. After the neutrons freeze-out, the unstable isotopes on the neutron-rich side will proceed to the stable isotopes mainly via $\beta$ decays, and the final abundances are obtained.
III Pseudo DRHBc mass table
The even-even and even-$Z$ odd-$N$ parts of the DRHBc mass table have been completed, with the numerical details and results summarized in Refs. [27, 77], respectively. Since the odd-$Z$ part of the DRHBc mass table is not yet available, a proper estimation for the masses of odd-$Z$ nuclei would provide preliminary insights into the knowledge of the whole nuclear landscape and the applications in the $r$-process simulation.
III.1 Mass estimation for odd nuclei
The mass of an odd nucleus can be estimated by the interpolation based on the properties of its neighboring even nuclei [93, 94]. Here we first give a brief introduction of the estimation, and then perform examinations by taking all even-$Z$ odd-$N$ nuclei in the nuclear chart with $8\leqslant Z\leqslant 120$ as examples.
The binding energy $E_{b}$ of a nucleus can be expressed in terms of the combinations of its neighbors’ quantities, for example,
@@ III.1 Mass estimation for odd nuclei
$-\frac{1}{2}[\delta_{n}(Z,N+1)+\delta_{n}(Z,N-1)],$ (12)
or $\text{or }E_{B}(Z,N)=$ $\frac{1}{2}[E_{b}(Z+1,N)+E_{b}(Z-1,N)]$
$-\frac{1}{2}[\delta_{p}(Z+1,N)+\delta_{p}(Z-1,N)],$ (13)
@@
$\delta_{n}(Z,N)$ $=\frac{(-1)^{N}}{2}[2E_{b}(Z,N)-E_{b}(Z,N+1)-E_{b}(Z,N-1)],$ (14)
$\delta_{p}(Z,N)$ $=\frac{(-1)^{Z}}{2}[2E_{b}(Z,N)-E_{b}(Z+1,N)-E_{b}(Z-1,N)].$ (15)
The odd-even mass difference is often approximated by the average pairing gap in Eq. (8) [43, 95, 96, 97],
@@
With this approximation, the $E_{b}$ of an odd nucleus can be obtained based on the $E_{b}$ and $\Delta_{n/p}$ of its neighboring even nuclei, where the latter ones have already been provided in a self-consistent and microscopic manner from the available part of the DRHBc mass table [27, 77]. Such an interpolation treatment on odd nuclei is simple, but contains the nuclear structure information from microscopic calculations to a certain extent, and thus ensures that the magnitude of pairing correlations is correct [93].
@@
To examine this interpolation treatment, taking all even-$Z$ odd-$N$ (even-odd for short) nuclei as examples and employing the above interpolation, the binding energies are obtained. Since these interpolated results are not from real DRHBc calculations, for convenience, they are labeled “pseudo” DRHBc results in this work. Figure 1 shows the differences between the pseudo binding energies and the real DRHBc results from self-consistent blocking calculations (labeled as “real”). In Fig. 1(a), where the rotational correction is not considered, for most nuclei the differences between the pseudo and real DRHBc results are less than 0.5 MeV, with the root-mean-square (rms) difference being 0.24 MeV, showing excellent agreement between these two treatments. In comparison, the rms deviation of real DRHBc binding energies (w/o rotational correction) from experimental data for even-$Z$ nuclei is 2.56 MeV [77]. The difference between the two theoretical treatments is dramatically smaller than the deviation from the data. This indicates that the pseudo DRHBc binding energies can well reproduce the real DRHBc results.
@@
For the density functional PC-PK1 adopted here, the rotational correction has been shown to play an important role in improving the binding energy description for deformed nuclei [98, 75, 76]. Therefore, it is necessary to compare $E_{\text{b}}^{\text{pseudo}}$ and $E_{\text{b}}^{\text{real}}$ after incorporating the rotational correction energy $E_{\text{rot}}$, and the corresponding results are shown in Fig. 1(b). The differences between these two treatments are still less than 0.5 MeV for most nuclei, while we also noticed that some nuclei exhibit larger differences than those in Fig. 1(a). For example, for the nuclei with $N=121,123$, $Z\approx 70$, as well as those with $N=77,79$, $Z\approx 67$, their $E_{\text{b+rot}}^{\text{pseudo}}-E_{\text{b+rot}}^{\text{real}}$ are larger than 1.5 MeV. These larger differences mainly correspond to the abrupt changes of nuclear shape, especially those near magic numbers, whose deformations suddenly decrease to (near-)zero. In Refs. [27, 77], it was mentioned that the cranking approximation used in DRHBc to obtain the $E_{\text{rot}}$ is not suitable for (near-)spherical nuclei, and therefore, for the nuclei with $|\beta_{2}|<0.05$, their $E_{\text{rot}}$ are taken as zero in the DRHBc mass table. Considering that the $E_{\text{b+rot}}^{\text{pseudo}}$ for an even-odd nucleus in Fig. 1(b) is interpolated based on the $E_{\text{b+rot}}^{\text{real}}$ of its neighboring even-even nuclei, the pseudo result may significantly deviate from the real result when the $\beta_{2}$ values of neighboring even-even nuclei straddle 0.05. Although some nuclei show larger differences, the rms difference in Fig. 1(b) is 0.49 MeV, which is slightly larger than that without $E_{\text{rot}}$ in Fig. 1(a), but still much smaller than the rms deviation of $E_{\text{b+rot}}^{\text{real}}$ from the data, 1.43 MeV. Therefore, after including the rotational correction, the pseudo and real DRHBc binding energies are still in good agreement for most nuclei, and their differences are not expected to substantially influence the discussions on physics.
III.2 Precision of the pseudo mass table
@@
Based on the DRHBc mass table for even-$Z$ nuclei [77] and using Eqs. (12) and (13) ($\delta_{n/p}$ are approximated by $\Delta_{n/p}$), the binding energies of odd-$Z$ nuclei with $8\leqslant Z\leqslant 120$ are estimated. Combining the pseudo binding energies of odd-$Z$ nuclei and the real DRHBc results of even-$Z$ nuclei available, a pseudo DRHBc mass table is obtained. For convenience, here the corresponding binding energy value is labeled as $E_{\text{b}}^{\text{pseudo}}$, which includes both pseudo results for odd-$Z$ nuclei and real DRHBc results for even-$Z$ nuclei. 9480 bound nuclei are obtained in total, where 2584 (27.3%) are even-even, 2245 (23.7%) are even-odd, 2513 (26.5%) are odd-even, and 2138 (22.6%) are odd-odd.
@@
Figure 2: The differences between the pseudo DRHBc binding energies and the available data [8] for the nuclei with $8\leqslant Z\leqslant 120$ scaled by colors. The details about the pseudo DRHBc results can be found in the text. (a) The rotational correction is not considered. (b) The rotational correction is included.
@@
Among these 9480 bound nuclei, the masses of 2413 nuclei have been measured experimentally [8]. Figure 2 shows the binding energy deviations $E_{b}^{\text{exp}}-E_{b}^{\text{pseudo}}$ for these measured nuclei. The results without $E_{\text{rot}}$ and with $E_{\text{rot}}$ are shown in panels (a) and (b), respectively. In Fig. 2(a), the deviation is relatively small near magic numbers, while it becomes larger when getting far away from magic numbers, which is related to the increase of deformation. The rms deviation of predicted binding energies from data is $\sigma=2.55$ MeV. In Fig. 2(b), after including the $E_{\text{rot}}$, the binding energy description is significantly improved for most nuclei, especially those with both neutron and proton numbers far away from magic numbers. Accordingly, the rms deviation reduces to $\sigma=1.47$ MeV. As a comparison, the corresponding rms deviation for real DRHBc binding energies for only even-$Z$ nuclei is $\sigma=2.56$ MeV without $E_{\text{rot}}$, and $\sigma=1.43$ MeV with $E_{\text{rot}}$ [77]. The rms deviations in Fig. 2 are very close to these two values, respectively, indicating that our pseudo binding energies of odd-$Z$ nuclei reach almost the same accuracy as the real DRHBc results of even-$Z$ nuclei.
@@
Figure 3: The differences between the pseudo DRHBc results of one-neutron separation energy and the available data [8] for the nuclei with $8\leqslant Z\leqslant 120$ scaled by colors. (a) The rotational correction is not considered. (b) The rotational correction is included.
@@
is intricately connected to the characteristics of the astrophysical $r$-process. Utilizing the $E_{b}^{\text{pseudo}}$ in Fig. 2(a), the $S_{n}^{\text{pseudo}}$ without rotational correction is calculated and its deviation from 2299 available data [8] is shown in Fig. 3(a). The deviations for most nuclei are within 1 MeV, and the rms deviation $\sigma=0.67$ MeV. Figure 3(b) shows the deviation of $S_{n}$ after including the rotational correction energy. It is found that, generally, $E_{\text{rot}}$ improves the description of $S_{n}$ at $Z>50$, but for light nuclei the deviation is increased, resulting in the rms deviation $\sigma=0.74$ MeV, slightly larger than the case without rotational correction in panel (a). The result that $E_{\text{rot}}$ does not improve the $\sigma$ of $S_{n}$ has been noticed in the calculations on even-$Z$ nuclei [77], and is attributed to the limit of cranking approximation, which has already been discussed in Fig. 1(b). We also noticed that the deviations in some odd-$Z$ nuclei are different from those of their even-$Z$ neighbors, forming odd-even staggering with respect to proton number in Fig. 3. This means that these pseudo results do not fully captures the odd-even effects. We expect that a fully self-consistent treatment in the complete DRHBc mass table in the future would eliminate this staggering in the deviation of $S_{n}$.
Table 1: The rms deviations $\sigma$ for binding energies ($E_{b}$), one- and two-neutron separation energies ($S_{n},S_{2n}$), as well as one- and two-proton separation energies ($S_{p},S_{2p}$) in the pseudo DRHBc mass table with respect to the available data [8] in the unit of MeV. For comparison, the corresponding deviations for even-$Z$ and odd-$Z$ nuclei are also listed separately.
@@
Similar to the discussions on Figs. 2 and 3, the rms deviations of other predicted quantities, including the two-neutron separation energy $S_{2n}$ and one- and two-proton separation energies $S_{p},S_{2p}$, in the pseudo DRHBc mass table are calculated and summarized in Table 1. It is noted that by introducing $E_{\text{rot}}$, the description of $E_{b}$ is significantly improved, while for separation energies the corresponding $\sigma$ values slightly increase, due to the cranking approximation as mentioned in the above discussions. For comparison, the corresponding results for even-$Z$ and odd-$Z$ nuclei are also listed separately in Table 1. It is found that odd-$Z$ nuclei have similar $\sigma$ values to even-$Z$ ones for both binding energies and separation energies, showing the consistency between the pseudo and real DRHBc results.
@@
In conclusion, in this Section, based on the available DRHBc results [77] for even-$Z$ nuclei, the binding energies of odd-$Z$ nuclei are estimated. A complete pseudo DRHBc mass table is obtained for all nuclei with $8\leqslant Z\leqslant 120$, with the accuracies in describing nuclear masses and separation energies expected to be close to the real DRHBc results from self-consistent calculations.
IV $R$-process simulation

### v2

@@ 2.2 Classical $r$-process
In the astrophysical environments with high-temperature $T\gtrsim 1.5~\mathrm{GK}$ and high neutron density $n_{n}\gtrsim 10^{20}\mathrm{~cm}^{-3}$, the equilibrium between neutron capture and photodisintegration reactions can be achieved, and the abundance ratios of neighboring isotopes on an isotopic chain can be obtained by the Saha equation Cowan1991PR ; Qian2003PPNP ; Arnould2007PR
$\frac{Y(Z,A+1)}{Y(Z,A)}=$ $n_{n}\left(\frac{2\pi\hbar^{2}}{m_{u}kT}\right)^{3/2}\frac{G(Z,A+1)}{2G(Z,A)}$ (11)
$\times\exp\left(\frac{A+1}{A}\right)^{3/2}\left[\frac{S_{n}(Z,A+1)}{kT}\right],$
where $Y(Z,A),S_{n}(Z,A)$, and $G(Z,A)$ are, respectively, the abundance, one-neutron separation energy, and partition function of nuclide $(Z,A)$, and $\hbar,k$, and $m_{u}$ are the Planck constant, Boltzmann constant, and atomic mass unit, respectively. Note that the neutron separation energy $S_{n}$ deduced from nuclear masses appears in the exponential, suggesting the importance of nuclear masses in the equilibrium.
@@
$\frac{dY(Z)}{dt}=$ $Y(Z-1)\sum_{A}P(Z-1,A)\lambda_{\beta}^{Z-1,A}$ (12)
$-Y(Z)\sum_{A}P(Z,A)\lambda_{\beta}^{Z,A},$
where $\lambda_{\beta}^{Z,A}$ is the $\beta$ decay rate of the nucleus $(Z,A)$, $Y(Z)=\sum_{A}Y(Z,A)=\sum_{A}P(Z,A)Y(Z)$ is the total abundance of each isotopic chain, and $P(Z,A)$ is the individual population coefficients obtained from the equilibrium condition in Eq. (11). By using Eqs. (11) and (12), the abundance of each isotope can be determined. After the neutrons freeze-out, the unstable isotopes on the neutron-rich side will proceed to the stable isotopes mainly via $\beta$ decays, and the final abundances are obtained.
3 Pseudo DRHBc mass table
The even-even and even-$Z$ odd-NN parts of the DRHBc mass table have been completed, with the numerical details and results summarized in Refs. Zhang2022ADNDT ; Guo2024ADNDT , respectively. Since the odd-$Z$ part of the DRHBc mass table is not yet available, a proper estimation for the masses of odd-$Z$ nuclei would provide preliminary insights into the knowledge of the whole nuclear landscape and the applications in the $r$-process simulation. In Section 3.1, the methodology of estimating the masses of odd nuclei based on the masses and pairing gaps of even nuclei is shown. In Section 3.2, the examination of the methodology is performed by comparing the estimated masses of even-$Z$ odd-NN (even-odd for short) nuclei with the corresponding results in the available DRHBc mass table, and the influence of different pairing gaps is discussed. In Section 3.3, by estimating the masses of odd-$Z$ nuclei based on the available DRHBc results of even-$Z$ nuclei, a pseudo DRHBc mass table is developed and its precision on mass descriptions is evaluated.
3.1 Mass estimation for odd nuclei
The mass of an odd nucleus can be estimated based on the properties of its neighboring even nuclei Kortelainen2010PRC ; Olsen2013PRL . The binding energy $E_{b}$ of a nucleus can be expressed in terms of the combinations of its neighbors’ quantities, for example,
@@ 3.1 Mass estimation for odd nuclei
$+\frac{(-1)^{N}}{2}[\delta_{n}(Z,N+1)+\delta_{n}(Z,N-1)],$ (13)
$\text{or }E_{b}(Z,N)=$ $\frac{1}{2}[E_{b}(Z+1,N)+E_{b}(Z-1,N)]$
$+\frac{(-1)^{Z}}{2}[\delta_{p}(Z+1,N)+\delta_{p}(Z-1,N)],$ (14)
@@
$\begin{aligned} \delta_{n}(Z,N)&=\frac{(-1)^{N}}{2}[2E_{b}(Z,N)-E_{b}(Z,N+1)\\ &\qquad-E_{b}(Z,N-1)],\end{aligned}$ (15)
$\begin{aligned} \delta_{p}(Z,N)&=\frac{(-1)^{Z}}{2}[2E_{b}(Z,N)-E_{b}(Z+1,N)\\ &\qquad-E_{b}(Z-1,N)].\end{aligned}$ (16)
The odd-even mass difference is often approximated by the microscopic average pairing gap in Eq. (8) Dobaczewski1984NPA ; Duguet2001PRC_2 ,
@@
With this approximation, the $E_{b}$ of an odd nucleus can be obtained based on the $E_{b}$ and $\Delta_{n/p}$ of its neighboring even nuclei, where the latter ones have already been provided in a self-consistent and microscopic manner from the available part of the DRHBc mass table Zhang2022ADNDT ; Guo2024ADNDT . Such an interpolation treatment on odd nuclei is simple, but contains the nuclear structure information from microscopic calculations to a certain extent, and thus ensures that the magnitude of pairing correlations is correct Kortelainen2010PRC .
3.2 Examination of the mass estimation: From even-even to even-odd nuclei
Figure 1: Average pairing gaps of even-even nuclei with $8\leqslant Z\leqslant 120$ in DRHBc theory for (a) neutron and (b) proton, in comparison with the empirical pairing gaps $\Delta=12A^{-1/2}$ and $\Delta=34A^{-3/4}$ Ring1980NMBP ; Bohr1998book , respectively.
In Fig. 1, the average pairing gaps calculated by Eq. (8) for the even-even nuclei with $8\leqslant Z\leqslant 120$ in DRHBc mass table are shown for neutron and proton, respectively. The empirical pairing gaps, $\Delta=12A^{-1/2}$ from the experimental odd-even mass staggering, and $\Delta=34A^{-3/4}$ from the semi-empirical mass formula of the liquid drop model Ring1980NMBP ; Bohr1998book , are also shown for comparison.
For the neutron pairing gap $\Delta_{n}$ in Fig. 1(a), generally, the DRHBc result decreases with the increase of $A$. The DRHBc results are in rough agreement with both empirical curves $12A^{-1/2}$ and $34A^{-3/4}$. Several peaks occur with the proton numbers at shell closures $Z=20,50$ and $82$, or at the pseudo shell $92$ Geng2006CPL ; Zhang2022ADNDT , while their neutron numbers mainly locate in the middle of the shells. This is attributed to the spherical shape at the proton shell closure and the high single-neutron level density in the middle of the neutron shell, leading to the strong pairing effect for neutron.
For the proton pairing gap $\Delta_{p}$ in Fig. 1(b), the DRHBc result shows a decreasing trend at $A<200$, while no significant decreasing or increasing trend occurs at $A>200$. This is because at a large $A$, the increase of $A$ mainly comes from the increase of $N$, instead of $Z$. Several peaks occur with the neutron numbers at shell closures $N=82,126,184$ and $258$, similar to the behavior of $\Delta_{n}$ in Fig. 1(a). Comparing Figs. 1 (a) and (b), it is found that $\Delta_{p}$ tends to locate above the empirical curves, while $\Delta_{n}$ tends to locate below the empirical ones at large $A$. Such difference originates from the general trend in the DRHBc calculations that pairing gap weakens as the number of the corresponding nucleon increases, i.e., $\Delta_{n}$ mainly depends on $N$ and $\Delta_{p}$ mainly depends on $Z$. Since $N>Z$ in most nuclei, as $A$ increases, $N$ grows faster than $Z$ in a general trend. This leads to a more rapid decreasing trend in $\Delta_{n}$ than in $\Delta_{p}$ for the DRHBc results and results in $\Delta_{n}$ lying below the empirical curves at large $A$, as the empirical ones only depend on $A$.
@@ 3.2 Examination of the mass estimation: From even-even to even-odd nuclei
To validate the methodology in Section 3.1, by taking all the even-odd nuclei in the nuclear chart with $8\leqslant Z\leqslant 120$ as examples, their binding energies $E_{b}$ are estimated based on the $E_{b}$ and $\Delta_{n}$ of the even-even nuclei, and compared with the available results in the DRHBc mass table Guo2024ADNDT . Since the estimated results are not obtained from the real DRHBc calculations, for convenience, they are labeled “pseudo” DRHBc results in this work. Figure 2 shows the differences between the pseudo binding energies and the real DRHBc results from self-consistent blocking calculations (labeled as “real”) for even-odd nuclei. In Fig. 2(a), where the rotational correction is not considered, for most nuclei the differences between the pseudo and real DRHBc results are less than 0.5 MeV, with the root-mean-square (rms) difference being 0.24 MeV, showing excellent agreement between these two treatments. In comparison, the rms deviation of real DRHBc binding energies (w/o rotational correction) from experimental data for even-$Z$ nuclei is 2.56 MeV Guo2024ADNDT . The difference between the two theoretical treatments is dramatically smaller than the deviation from the data. This indicates that the pseudo DRHBc binding energies can well reproduce the real DRHBc results.
@@
For the density functional PC-PK1 adopted here, the rotational correction has been shown to play an important role in improving the binding energy description for deformed nuclei Zhao2010PRC ; Zhang2020PRC ; Pan2022PRC . Therefore, it is necessary to compare $E_{\text{b}}^{\text{pseudo}}$ and $E_{\text{b}}^{\text{real}}$ after incorporating the rotational correction energy $E_{\text{rot}}$, and the corresponding results are shown in Fig. 2(b). The differences between these two treatments are still less than 0.5 MeV for most nuclei, while we also noticed that some nuclei exhibit larger differences than those in Fig. 2(a). For example, for the nuclei with $N=121,123$, $Z\approx 70$, as well as those with $N=77,79$, $Z\approx 67$, their $E_{\text{b+rot}}^{\text{pseudo}}-E_{\text{b+rot}}^{\text{real}}$ are larger than 1.5 MeV. These larger differences mainly correspond to the abrupt changes of nuclear shape, especially those near magic numbers, whose deformations suddenly decrease to (near-)zero. In this case, the cranking approximation is not suitable for $E_{\text{rot}}$ in Eq. (9). Considering that the $E_{\text{b+rot}}^{\text{pseudo}}$ for an even-odd nucleus in Fig. 2(b) is interpolated based on the $E_{\text{b+rot}}^{\text{real}}$ of its neighboring even-even nuclei, the pseudo result may significantly deviate from the real result when the $\beta_{2}$ values of neighboring even-even nuclei straddle 0.05. Although some nuclei show larger differences, the rms difference in Fig. 2(b) is 0.41 MeV, which is slightly larger than that without $E_{\text{rot}}$ in Fig. 2(a), but still much smaller than the rms deviation of $E_{\text{b+rot}}^{\text{real}}$ from the data, 1.43 MeV. Therefore, after including the rotational correction, the pseudo and real DRHBc binding energies are still in good agreement for most nuclei, and their differences are not expected to substantially influence the discussions on physics.
To further examine the methodology, based on the masses of even-even nuclei and the empirical pairing gaps $\Delta=12A^{-1/2}$ and $\Delta=34A^{-3/4}$, the binding energies of even-odd nuclei are also estimated, respectively, for comparisons. In Fig. 3, the estimated binding energies are compared with the real DRHBc results.
Figure 3: The same as Fig. 2, but the pairing gaps are estimated with the empirical formulas (a) $\Delta=12A^{-1/2}$ (“emp-1/2” for short) and (b) $\Delta=34A^{-3/4}$ (“emp-3/4” for short). The rotational correction is included.
For the estimation based on the empirical pairing gap $\Delta=12A^{-1/2}$ shown in Fig. 3(a), the rms difference is $\sigma=0.46$ MeV, larger than that in Fig. 2(b). For the nuclei with $Z<28$, noticeable differences with $E_{\text{b+rot}}^{\text{emp-1/2}}-E_{\text{b+rot}}^{\text{real}}\approx-2$ MeV are observed. Such differences can be understood from Fig. 1(a), where the $12A^{-1/2}$ curve is higher than most of the DRHBc results when $A$ is small, leading the underestimations of the binding energies according to Eq. (13).
For the estimation based on the empirical pairing gap $\Delta=34A^{-3/4}$ shown in Fig. 3(b), the rms difference is $\sigma=0.49$ MeV, larger than those in Fig. 2(b) and 3(a). For the nuclei with $Z<28$, the differences reach $E_{\text{b+rot}}^{\text{emp-3/4}}-E_{\text{b+rot}}^{\text{real}}\approx-3$ MeV, larger than those in Fig. 3(a). This is attributed to the fact that the $34A^{-1/2}$ curve is higher than the $12A^{-1/2}$ curve when $A$ is small in Fig. 1(a), leading to more underestimated binding energies according to Eq. (13).
According to above discussions, the pseudo DRHBc results nicely reproduce the binding energies in the real DRHBc mass table for even-odd nuclei. The comparisons with the results based on the empirical pairing gaps show that the global behavior of the binding energy is not very sensitive to the details in the evolution of pairing gaps.
3.3 The pseudo mass table: From even-$Z$ to odd-$Z$ nuclei
@@
Based on the DRHBc mass table for even-$Z$ nuclei Guo2024ADNDT and using Eqs. (13) and (14) ($\delta_{n/p}$ are approximated by $\Delta_{n/p}$), the binding energies of odd-$Z$ nuclei with $8\leqslant Z\leqslant 120$ are estimated. Combining the pseudo binding energies of odd-$Z$ nuclei and the real DRHBc results of even-$Z$ nuclei available, a pseudo DRHBc mass table is obtained. For convenience, here the corresponding binding energy value is labeled as $E_{\text{b}}^{\text{pseudo}}$, which includes both pseudo results for odd-$Z$ nuclei and real DRHBc results for even-$Z$ nuclei. 9480 bound nuclei are obtained in total, where 2584 (27.3%) are even-even, 2245 (23.7%) are even-odd, 2513 (26.5%) are odd-even, and 2138 (22.6%) are odd-odd.
@@
Figure 4: The differences between the pseudo DRHBc binding energies and the available data Wang2021CPC for the nuclei with $8\leqslant Z\leqslant 120$ scaled by colors. The details about the pseudo DRHBc results can be found in the text. (a) The rotational correction is not considered. (b) The rotational correction is included.
@@
Among these 9480 bound nuclei, the masses of 2413 nuclei have been measured experimentally Wang2021CPC . Figure 4 shows the binding energy deviations $E_{b}^{\text{exp}}-E_{b}^{\text{pseudo}}$ for these measured nuclei. The results without $E_{\text{rot}}$ and with $E_{\text{rot}}$ are shown in panels (a) and (b), respectively. In Fig. 4(a), the deviation is relatively small near magic numbers, while it becomes larger when getting far away from magic numbers, which is related to the increase of deformation. The rms deviation of predicted binding energies from data is $\sigma=2.55$ MeV. In Fig. 4(b), after including the $E_{\text{rot}}$, the binding energy description is significantly improved for most nuclei, especially those with both neutron and proton numbers far away from magic numbers. Accordingly, the rms deviation reduces to $\sigma=1.47$ MeV. As a comparison, the corresponding rms deviation for real DRHBc binding energies for only even-$Z$ nuclei is $\sigma=2.56$ MeV without $E_{\text{rot}}$, and $\sigma=1.43$ MeV with $E_{\text{rot}}$ Guo2024ADNDT . The rms deviations in Fig. 4 are very close to these two values, respectively, indicating that our pseudo binding energies of odd-$Z$ nuclei reach almost the same accuracy as the real DRHBc results of even-$Z$ nuclei.
@@
Figure 5: The differences between the pseudo DRHBc results of one-neutron separation energy and the available data Wang2021CPC for the nuclei with $8\leqslant Z\leqslant 120$ scaled by colors. (a) The rotational correction is not considered. (b) The rotational correction is included.
@@
is intricately connected to the characteristics of the astrophysical $r$-process. Utilizing the $E_{b}^{\text{pseudo}}$ in Fig. 4(a), the $S_{n}^{\text{pseudo}}$ without rotational correction is calculated and its deviation from 2299 available data Wang2021CPC is shown in Fig. 5(a). The deviations for most nuclei are within 1 MeV, and the rms deviation $\sigma=0.67$ MeV. Figure 5(b) shows the deviation of $S_{n}$ after including the rotational correction energy. It is found that, generally, $E_{\text{rot}}$ improves the description of $S_{n}$ at $Z>50$, but for light nuclei the deviation is increased, resulting in the rms deviation $\sigma=0.74$ MeV, slightly larger than the case without rotational correction in panel (a). The result that $E_{\text{rot}}$ does not improve the $\sigma$ of $S_{n}$ has been noticed in the calculations on even-$Z$ nuclei Guo2024ADNDT , and is attributed to the limit of cranking approximation, which has already been discussed in Fig. 2(b). We also noticed that the deviations in some odd-$Z$ nuclei are different from those of their even-$Z$ neighbors, forming odd-even staggering with respect to proton number in Fig. 5. This means that these pseudo results do not fully captures the odd-even effects. We expect that a fully self-consistent treatment in the complete DRHBc mass table in the future would eliminate this staggering in the deviation of $S_{n}$.
Table 1: The rms deviations $\sigma$ for binding energies ($E_{b}$), one- and two-neutron separation energies ($S_{n},S_{2n}$), as well as one- and two-proton separation energies ($S_{p},S_{2p}$) in the pseudo DRHBc mass table with respect to the available data Wang2021CPC in the unit of MeV. For comparison, the corresponding deviations for even-$Z$ and odd-$Z$ nuclei are also listed separately.
@@
Similar to the discussions on Figs. 4 and 5, the rms deviations of other predicted quantities, including the two-neutron separation energy $S_{2n}$ and one- and two-proton separation energies $S_{p},S_{2p}$, in the pseudo DRHBc mass table are calculated and summarized in Table 1. It is noted that by introducing $E_{\text{rot}}$, the description of $E_{b}$ is significantly improved, while for separation energies the corresponding $\sigma$ values slightly increase, due to the cranking approximation as mentioned in the above discussions. For comparison, the corresponding results for even-$Z$ and odd-$Z$ nuclei are also listed separately in Table 1. It is found that odd-$Z$ nuclei have similar $\sigma$ values to even-$Z$ ones for both binding energies and separation energies, showing the consistency between the pseudo and real DRHBc results.
In addition, based on the empirical pairing gaps $\Delta=12A^{-1/2}$ and $\Delta=34A^{-3/4}$, the masses of odd-$Z$ nuclei are estimated in the same manner. The corresponding $\sigma(E_{b})$ (w/o $E_{\mathrm{rot}}$) are 2.66 MeV and 2.61 MeV, respectively. The corresponding $\sigma(E_{b})$ (w/ $E_{\mathrm{rot}}$) are 1.47 MeV and 1.43 MeV, respectively. These values are close to the $\sigma(E_{b})$ in Table 1. The impacts on $\sigma(E_{b})$ due to the differences among these three sets of pairing gaps are less than 0.1 MeV. Such small differences are consistent with the conclusion in Section 3.2 that the global behavior of $E_{b}$ is not very sensitive to the details in the evolution of pairing gaps.
@@
In conclusion, in this Section, based on the available DRHBc results Guo2024ADNDT for even-$Z$ nuclei, the binding energies of odd-$Z$ nuclei are estimated. A complete pseudo DRHBc mass table is obtained for all nuclei with $8\leqslant Z\leqslant 120$, with the accuracies in describing nuclear masses and separation energies expected to be close to the real DRHBc results from self-consistent calculations.
4 $R$-process simulation